The sum of the two digits of a two-digit number is 12.If the digits are interchanged,the number is increased by 36. Find the number
Answer by Alan3354(69443)

The sum of the two digits of a two-digit number is 12.If the digits are interchanged,the number is increased by 36. Find the number
------------
T + U = 12
T - U = 36/9 = 4
-----------------
If that doesn't work, reverse the T and U digits.